OpenStudy (anonymous):

how would you do the problem log5.4-log0.054?

OpenStudy (anonymous):

log5.4-log0.054 // using the following log rule log(x)-log(y) = log(x/y) log(5.4)-log(.054) = log(5.4/.054) = ?